Required: Use your knowledge of the financial statement equations and their interrelationships to fill in the missing amounts ota Revenues Tota Expenses Net Income Loss Tota Assets Tota Liabilities Stockholders Equity Company Vanilla Co Chocolate, Inc. Strawberry Co 20,100 (29,600) 60,200 33,200 66,200 38,200 101,000 130,400 60,200 30,200 10,100 40,100
Answer #1

Net income=Revenue-Expenses

Total assets=Total liabilities+Stockholders equity

Company Total revenue Total expenses Net income Total assets Total liabilities Stockholders equity
Vanilla Co. 60200 (60200-20100)=$40100 20100 (33200+66200)=$99400 33200 66200
Chocolate Inc (101000-29600)=$71400 101000 (29600) 130400 (130400-38200)=$92200 38200
Strawberry Co. 30200 10100 (30200-10100)=$20100 60200 40100 (60200-40100)=$20100